# Arithmetic of cuts in ordered abelian groups and of ideals over valuation rings

### v4

@@ 2.5. Partial answers to questions (QFS1), (QFS2) and (QFS3)
so equality holds everywhere. It follows that $S_{2}+\alpha^{\prime}-\alpha=S_{1}+\alpha+\alpha^{\prime}-\alpha=S_{1}+\alpha^{\prime}=S_{2}\,$, whence $\alpha^{\prime}-\alpha\in{\mathcal{G}}(S_{2})=\{0\}$. Thus $\alpha^{\prime}=\tau$, a contradiction. ∎
@@
As stated in the introduction, $T=S_{2}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T\subseteq S_{2}$; hence if a solution of equation (1) exists, then $T=S_{2}-_{\rm fs}S_{1}$ is the largest solution of it. For $S_{1}=\alpha^{-}$, our assertion follows from Lemma 2.5. If $S_{1}$ is nonprincipal, then by assumption, $S_{2}$ is nonprincipal, and it follows from part 2) of Corollary 2.19 that $S_{2}-S_{1}^{c}$ is a solution of (1). Hence our statement for this case follows from Lemma 2.21. ∎
@@ 2.7. The equation $S_{1}+T=S_{2}$ in ordered abelian groups of arbitrary rank
1) The preimage of $-(T/{\mathcal{G}}(T))^{c}$ under $\varphi_{{\mathcal{G}}(T)}$ is $-T^{c}$.
2) For each $\alpha\in G$, the preimage of $(\alpha/{\mathcal{G}}(T))^{-}$ under $\varphi_{{\mathcal{G}}(T)}$ is $a^{-}+{\mathcal{G}}(T)$.
@@
1): By part 1) of Lemma 2.16, ${\mathcal{G}}(-T^{c})={\mathcal{G}}(T^{c})={\mathcal{G}}(T)$, hence the preimage of $-T^{c}/{\mathcal{G}}(T)$ under $\varphi_{{\mathcal{G}}(T)}$ is $-T^{c}+{\mathcal{G}}(T)=-T^{c}$. It remains to show that $-T^{c}/{\mathcal{G}}(T)=-(T/{\mathcal{G}}(T))^{c}$. Since ${\mathcal{G}}(T^{c})={\mathcal{G}}(T)\,$, the sets $T/{\mathcal{G}}(T)$ and $T^{c}/{\mathcal{G}}(T)$ are disjoint. On the other hand, their union is $G/{\mathcal{G}}(T)\,$, so $T^{c}/{\mathcal{G}}(T)=(T/{\mathcal{G}}(T))^{c}$, showing that $-T^{c}/{\mathcal{G}}(T)=-(T/{\mathcal{G}}(T))^{c}$ holds.
@@
()(*)( ∗ )  ${\mathcal{G}}(S_{1})\subseteq{\mathcal{G}}(S_{2})$, and $S_{2}/{\mathcal{G}}(S_{2})$ is nonprincipal if $S_{1}/{\mathcal{G}}(S_{2})$ is nonprincipal
@@
If condition ()(*)( ∗ ) does not hold, or equivalently, one of the cases
@@
To simplify notation, we set ${\mathcal{G}}_{2}:={\mathcal{G}}(S_{2})$. Assume that condition ()(*)( ∗ ) holds. From part 2) of Proposition 2.25 we know that ${\mathcal{G}}(S_{1}/{\mathcal{G}}_{2})={\mathcal{G}}(S_{2}/{\mathcal{G}}_{2})=\{0\}$. Hence by Proposition 2.23, the largest solution of $S_{1}/{\mathcal{G}}_{1}+\overline{T}=S_{2}/{\mathcal{G}}_{1}$ is $\overline{T}=S_{2}/{\mathcal{G}}_{2}-\alpha/{\mathcal{G}}_{2}=(S_{2}-\alpha)/{\mathcal{G}}_{2}$ if $S_{1}/{\mathcal{G}}_{2}=(\alpha/{\mathcal{G}}_{2})^{-}$, and $\overline{T}=(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})\,\widehat{}$ if $S_{1}/{\mathcal{G}}_{2}$ is nonprincipal.
@@
If $\overline{T}=(S_{2}-\alpha)/{\mathcal{G}}_{2}\,$, then $T=(S_{2}-\alpha)+{\mathcal{G}}_{2}=S_{2}-\alpha$ by part 3) of Lemma 2.13. Now assume that $S_{1}/{\mathcal{G}}_{2}$ is nonprincipal and $\overline{T}=(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})\,\widehat{}$. By part 4) of Proposition 2.25, the former implies that ${\mathcal{G}}_{2}\subseteq{\mathcal{G}}(S_{1})$, whence ${\mathcal{G}}(S_{1})={\mathcal{G}}_{2}\,$. Therefore, $\varphi_{{\mathcal{G}}_{2}}^{-1}(-(S_{1}/{\mathcal{G}}_{2})^{c})=-S_{1}^{c}$ by part 1) of Lemma 2.29. By Lemma 2.24, $\varphi_{{\mathcal{G}}_{2}}^{-1}(S_{2}/{\mathcal{G}}_{2})=S_{2}+{\mathcal{G}}_{2}=S_{2}\,$, hence $\varphi_{{\mathcal{G}}_{2}}^{-1}(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})=S_{2}-S_{1}^{c}$. Thus the preimage of $\overline{T}$ is $T=(S_{2}-S_{1}^{c})^{\diamondsuit}$.
Now assume that condition ()(*)( ∗ ) does not hold. Then one of the cases i) or ii) must hold; observe that if ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}_{2}$, then by part 4) of Proposition 2.25, $S_{1}/{\mathcal{G}}_{2}$ is principal and condition ()(*)( ∗ ) holds. Note that in both cases, $S_{1}$ is nonprincipal. In case i), Proposition 2.15 shows that equation (1) has no solution. Assume that case ii) holds and that the final segment $T$ is a solution of equation (1). Then $S_{1}/{\mathcal{G}}_{2}+T/{\mathcal{G}}_{2}=S_{2}/{\mathcal{G}}_{2}$, but this contradicts part 2) of Lemma 2.4 applied to the group $G/{\mathcal{G}}_{2}\,$. ∎
@@
2) Assume that ${\mathcal{G}}(S_{1})={\mathcal{G}}(S_{2})$ and $S_{1}/{\mathcal{G}}(S_{1})$ is nonprincipal, but $S_{2}/{\mathcal{G}}(S_{1})$ is principal, say $S_{2}/{\mathcal{G}}(S_{1})=(\alpha/{\mathcal{G}}(S_{1}))^{-}$ for some $\alpha\in S_{2}\,$. Then $S^{\prime}_{2}:=(\alpha+{\mathcal{G}}(S_{1}))^{+}\subsetneq S_{2}$ is the largest final segment contained in $S_{2}$ that is nonprincipal.
In both cases, condition ()(*)( ∗ ) of Theorem 2.33 is satisfied with $S^{\prime}_{2}$ in place of $S_{2}$ and $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$. Further, $S_{2}^{\prime}-_{\rm fs}S_{1}=S_{2}-_{\rm fs}S_{1}$.
@@
We observe that $S^{\prime}_{2}$ is the largest final segment contained in $S_{2}$ having invariance group ${\mathcal{G}}_{1}$. Indeed, otherwise there would exist an element $\beta\in(\alpha+{\mathcal{G}}_{1})\cap S_{2}$ such that $\beta+{\mathcal{G}}_{1}\subseteq S_{2}\,$. However, it would follow that $\beta-\alpha\in{\mathcal{G}}_{1}$ and $\alpha+{\mathcal{G}}_{1}=\beta+{\mathcal{G}}_{1}\subseteq S_{2}\,$, contradiction. Consequently, if $S_{1}+T\subseteq S_{2}$, then we must have that $S_{1}+T\subseteq S^{\prime}_{2}$ because of ${\mathcal{G}}_{1}\subseteq{\mathcal{G}}(S_{1}+T)$ by part 2) of Lemma 2.13.
@@
Assume that $S^{\prime}_{2}/{\mathcal{G}}_{1}=(\alpha/{\mathcal{G}}_{1})^{+}$ is principal; it then follows that $G/{\mathcal{G}}_{1}$ is discretely ordered. Hence by part 5) of Proposition 2.25, also $S_{1}/{\mathcal{G}}_{1}$ is principal. This shows that condition ()(*)( ∗ ) of Theorem 2.33 holds with $S_{2}$ replaced by $S_{2}^{\prime}\,$. Now it follows from Theorem 2.33 that $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$. Since $S_{2}^{\prime}\subset S_{2}$, we have $S_{2}^{\prime}-_{\rm fs}S_{1}\subseteq S_{2}-_{\rm fs}S_{1}\,$. On the other hand, ${\mathcal{G}}_{1}\subseteq{\mathcal{G}}(S_{1}+(S_{2}-_{\rm fs}S_{1}))$ which implies that $S_{1}+(S_{2}-_{\rm fs}S_{1})\subseteq S_{2}^{\prime}$ and therefore $S_{2}-_{\rm fs}S_{1}\subseteq S_{2}^{\prime}-_{\rm fs}S_{1}\,$.
@@
2): As in the proof of part 1) it is shown that $S^{\prime}_{2}\subsetneq S_{2}$ and that ${\mathcal{G}}(S^{\prime}_{2})={\mathcal{G}}_{1}$. Also as in the proof of part 1) it is shown that if $S^{\prime}_{2}/{\mathcal{G}}_{1}=(\alpha/{\mathcal{G}}_{1})^{+}$ is principal, then so is $S_{1}/{\mathcal{G}}_{1}$. But this contradicts our assumption, so $(\alpha/{\mathcal{G}}_{1})^{+}$ is nonprincipal. Since $\alpha+{\mathcal{G}}_{1}=\alpha+{\mathcal{G}}(S_{2})$ is an initial segment od $S_{2}\,$, any final segment $S$ that properly contains $S^{\prime}_{2}$ will contain an element $\beta\in\alpha+{\mathcal{G}}_{1}\,$. This implies that $S/{\mathcal{G}}_{1}$ contains $S_{2}/{\mathcal{G}}_{1}\,$, so if it is contained in $S_{2}/{\mathcal{G}}_{1}\,$, then it is equal to $S_{2}/{\mathcal{G}}_{1}$ and hence principal. This proves that $S^{\prime}_{2}$ is the largest final segment contained in $S_{2}$ such that $S^{\prime}_{2}/{\mathcal{G}}_{1}$ is nonprincipal.
@@
Again, condition ()(*)( ∗ ) of Theorem 2.33 holds with $S_{2}$ replaced by $S_{2}^{\prime}$ so that $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$ and as before, $S_{2}^{\prime}-_{\rm fs}S_{1}\subseteq S_{2}-_{\rm fs}S_{1}\,$. To show the reverse inclusion, suppose that $S_{2}^{\prime}\subsetneq S_{1}+(S_{2}-_{\rm fs}S_{1})$. Then $(S_{1}+(S_{2}-_{\rm fs}S_{1}))/{\mathcal{G}}_{1}$ contains $\alpha/{\mathcal{G}}_{1}$ and thus $S_{1}+(S_{2}-_{\rm fs}S_{1})$ must be equal to $S_{2}\,$, which is impossible by Theorem 2.33. Hence again, $S_{1}+(S_{2}-_{\rm fs}S_{1})\subseteq S_{2}^{\prime}$ and therefore $S_{2}-_{\rm fs}S_{1}\subseteq S_{2}^{\prime}-_{\rm fs}S_{1}\,$. ∎

### v5

@@ 2.5. Partial answers to questions (QFS1), (QFS2) and (QFS3)
so equality holds everywhere. It follows that $S_{2}+\alpha^{\prime}-\alpha=S_{1}+\alpha+\alpha^{\prime}-\alpha=S_{1}+\alpha^{\prime}=S_{2}\,$, whence $\alpha^{\prime}-\alpha\in{\mathcal{G}}(S_{2})=\{0\}$. Thus $\alpha^{\prime}=\alpha$, a contradiction. ∎
@@
As stated in the introduction, $T=S_{2}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T\subseteq S_{2}$; hence if a solution of equation (1) exists, then $T=S_{2}-_{\rm fs}S_{1}$ is the largest solution of it. For $S_{1}=\alpha^{-}$, our assertion follows from Proposition 2.5. If $S_{1}$ is nonprincipal, then by assumption, $S_{2}$ is nonprincipal, and it follows from part 2) of Corollary 2.19 that $S_{2}-S_{1}^{c}$ is a solution of (1). Hence our statement for this case follows from Lemma 2.21. ∎
@@ 2.7. The equation $S_{1}+T=S_{2}$ in ordered abelian groups of arbitrary rank
1) The preimage of $(T/{\mathcal{G}}(T))^{c}$ under $\varphi_{{\mathcal{G}}(T)}$ is $T^{c}$.
2) For each $\alpha\in G$, the preimage of $(\alpha/{\mathcal{G}}(T))^{-}$ under $\varphi_{{\mathcal{G}}(T)}$ is $\alpha^{-}+{\mathcal{G}}(T)$.
@@
1): By part 1) of Lemma 2.16, ${\mathcal{G}}(-T^{c})={\mathcal{G}}(T^{c})={\mathcal{G}}(T)$, hence the preimage of $T^{c}/{\mathcal{G}}(T)$ under $\varphi_{{\mathcal{G}}(T)}$ is $T^{c}+{\mathcal{G}}(T)=-T^{c}$. It remains to show that $T^{c}/{\mathcal{G}}(T)=-(T/{\mathcal{G}}(T))^{c}$. Since ${\mathcal{G}}(T^{c})={\mathcal{G}}(T)\,$, the sets $T/{\mathcal{G}}(T)$ and $T^{c}/{\mathcal{G}}(T)$ are disjoint. On the other hand, their union is $G/{\mathcal{G}}(T)\,$, so $T^{c}/{\mathcal{G}}(T)=(T/{\mathcal{G}}(T))^{c}$, showing that $T^{c}/{\mathcal{G}}(T)=-(T/{\mathcal{G}}(T))^{c}$ holds.
@@
()(*) ${\mathcal{G}}(S_{1})\subseteq{\mathcal{G}}(S_{2})$, and $S_{2}/{\mathcal{G}}(S_{2})$ is nonprincipal if $S_{1}/{\mathcal{G}}(S_{2})$ is nonprincipal
@@
If condition ()(*) does not hold, or equivalently, one of the cases
@@
To simplify notation, we set ${\mathcal{G}}_{2}:={\mathcal{G}}(S_{2})$. Assume that condition ()(*) holds. From part 2) of Proposition 2.25 we know that ${\mathcal{G}}(S_{1}/{\mathcal{G}}_{2})={\mathcal{G}}(S_{2}/{\mathcal{G}}_{2})=\{0\}$. Hence by Proposition 2.23, the largest solution of $S_{1}/{\mathcal{G}}_{2}+\overline{T}=S_{2}/{\mathcal{G}}_{2}$ is $\overline{T}=S_{2}/{\mathcal{G}}_{2}-\alpha/{\mathcal{G}}_{2}=(S_{2}-\alpha)/{\mathcal{G}}_{2}$ if $S_{1}/{\mathcal{G}}_{2}=(\alpha/{\mathcal{G}}_{2})^{-}$, and $\overline{T}=(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})\,\widehat{}\$if $S_{1}/{\mathcal{G}}_{2}$ is nonprincipal.
@@
If $\overline{T}=(S_{2}-\alpha)/{\mathcal{G}}_{2}\,$, then $T=(S_{2}-\alpha)+{\mathcal{G}}_{2}=S_{2}-\alpha$ by Lemma 2.24 and part 3) of Lemma 2.13. Now assume that $S_{1}/{\mathcal{G}}_{2}$ is nonprincipal and $\overline{T}=(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})\,\widehat{}\$. By part 4) of Proposition 2.25, the former implies that ${\mathcal{G}}_{2}\subseteq{\mathcal{G}}(S_{1})$, whence ${\mathcal{G}}(S_{1})={\mathcal{G}}_{2}\,$. Therefore, $\varphi_{{\mathcal{G}}_{2}}^{-1}(-(S_{1}/{\mathcal{G}}_{2})^{c})=-S_{1}^{c}$ by part 1) of Lemma 2.29. By Lemma 2.24, $\varphi_{{\mathcal{G}}_{2}}^{-1}(S_{2}/{\mathcal{G}}_{2})=S_{2}+{\mathcal{G}}_{2}=S_{2}\,$, hence $\varphi_{{\mathcal{G}}_{2}}^{-1}(S_{2}/{\mathcal{G}}_{2}-(S_{1}/{\mathcal{G}}_{2})^{c})=S_{2}-S_{1}^{c}$. Thus the preimage of $\overline{T}$ is $T=(S_{2}-S_{1}^{c})^{\diamondsuit}$.
Now assume that condition ()(*) does not hold. Then one of the cases i) or ii) must hold; observe that if ${\mathcal{G}}(S_{1})\subsetneq{\mathcal{G}}_{2}$, then by part 4) of Proposition 2.25, $S_{1}/{\mathcal{G}}_{2}$ is principal and condition ()(*) holds. Note that in both cases, $S_{1}$ is nonprincipal. In case i), Proposition 2.15 shows that equation (1) has no solution. Assume that case ii) holds and that the final segment $T$ is a solution of equation (1). Then $S_{1}/{\mathcal{G}}_{2}+T/{\mathcal{G}}_{2}=S_{2}/{\mathcal{G}}_{2}$, but this contradicts part 2) of Lemma 2.4 applied to the group $G/{\mathcal{G}}_{2}\,$. ∎
@@
2) Assume that ${\mathcal{G}}(S_{1})={\mathcal{G}}(S_{2})$ and $S_{1}/{\mathcal{G}}(S_{1})$ is nonprincipal, but $S_{2}/{\mathcal{G}}(S_{1})$ is principal, say $S_{2}/{\mathcal{G}}(S_{1})=(\alpha/{\mathcal{G}}(S_{1}))^{-}$ for some $\alpha\in S_{2}\,$. Then $S^{\prime}_{2}:=(\alpha+{\mathcal{G}}(S_{1}))^{+}\subsetneq S_{2}$ is the largest final segment contained in $S_{2}$ such that $S^{\prime}_{2}/{\mathcal{G}}(S_{1})$ is nonprincipal.
In both cases, condition ()(*) of Theorem 2.33 is satisfied with $S^{\prime}_{2}$ in place of $S_{2}$ and $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$. Further, $S_{2}^{\prime}-_{\rm fs}S_{1}=S_{2}-_{\rm fs}S_{1}$.
@@
Suppose that there is a final segment $S$ of $S_{2}$ properly containing $S^{\prime}_{2}$ and such that ${\mathcal{G}}_{1}\subseteq{\mathcal{G}}(S)$. Then there would exist an element $\beta\in(\alpha+{\mathcal{G}}_{1})\cap S_{2}$ such that $\beta+{\mathcal{G}}_{1}\subseteq\beta+{\mathcal{G}}(S)\subseteq S\subseteq S_{2}\,$. However, it would follow that $\beta-\alpha\in{\mathcal{G}}_{1}$ and $\alpha+{\mathcal{G}}_{1}=\beta+{\mathcal{G}}_{1}\subseteq S_{2}\,$, contradiction. This shows that $S^{\prime}_{2}$ is the largest final segment contained in $S_{2}$ having an invariance group containing ${\mathcal{G}}_{1}$.
Now if $S_{1}+T\subseteq S_{2}$, then ${\mathcal{G}}_{1}\subseteq{\mathcal{G}}(S_{1}+T)$ by part 2) of Lemma 2.13, hence by what we have just shown, $S_{1}+T\subseteq S^{\prime}_{2}\,$.
@@
Assume that $S^{\prime}_{2}/{\mathcal{G}}_{1}=(\alpha/{\mathcal{G}}_{1})^{+}$ is principal; it then follows that $G/{\mathcal{G}}_{1}$ is discretely ordered. Hence by part 5) of Proposition 2.25, also $S_{1}/{\mathcal{G}}_{1}$ is principal. This shows that condition ()(*) of Theorem 2.33 holds with $S_{2}$ replaced by $S_{2}^{\prime}\,$. Now it follows from Theorem 2.33 that $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$. Since $S_{2}^{\prime}\subset S_{2}$, we have $S_{2}^{\prime}-_{\rm fs}S_{1}\subseteq S_{2}-_{\rm fs}S_{1}\,$. On the other hand, ${\mathcal{G}}_{1}\subseteq{\mathcal{G}}(S_{1}+(S_{2}-_{\rm fs}S_{1}))$ which implies that $S_{1}+(S_{2}-_{\rm fs}S_{1})\subseteq S_{2}^{\prime}$ and therefore $S_{2}-_{\rm fs}S_{1}\subseteq S_{2}^{\prime}-_{\rm fs}S_{1}\,$.
@@
2): As in the proof of part 1) it is shown that $S^{\prime}_{2}\subsetneq S_{2}$ and that ${\mathcal{G}}(S^{\prime}_{2})={\mathcal{G}}_{1}$. Also as in the proof of part 1) it is shown that if $S^{\prime}_{2}/{\mathcal{G}}_{1}=(\alpha/{\mathcal{G}}_{1})^{+}$ is principal, then so is $S_{1}/{\mathcal{G}}_{1}$. But this contradicts our assumption, so $S^{\prime}_{2}/{\mathcal{G}}_{1}$ is nonprincipal.
Any final segment $S$ that properly contains $S^{\prime}_{2}=(\alpha+{\mathcal{G}}_{1})^{+}$ will contain an element $\beta\in\alpha+{\mathcal{G}}_{1}\,$. This implies that $S/{\mathcal{G}}_{1}$ contains $\alpha/{\mathcal{G}}_{1}$ and thus contains $(\alpha/{\mathcal{G}}(S_{1}))^{-}=S_{2}/{\mathcal{G}}_{1}\,$. Hence if $S$ is contained in $S_{2}\,$, $S/{\mathcal{G}}_{1}$ is contained in and therefore equal to $S_{2}/{\mathcal{G}}_{1}\,$, which is principal. This proves that $S^{\prime}_{2}$ is the largest final segment contained in $S_{2}$ such that $S^{\prime}_{2}/{\mathcal{G}}_{1}$ is nonprincipal.
@@
Again, condition ()(*) of Theorem 2.33 holds with $S_{2}$ replaced by $S_{2}^{\prime}$ so that $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$ and as before, $S_{2}^{\prime}-_{\rm fs}S_{1}\subseteq S_{2}-_{\rm fs}S_{1}\,$. To show the reverse inclusion, suppose that $S_{2}^{\prime}\subsetneq S_{1}+(S_{2}-_{\rm fs}S_{1})$. Then $(S_{1}+(S_{2}-_{\rm fs}S_{1}))/{\mathcal{G}}_{1}$ contains $\alpha/{\mathcal{G}}_{1}$ and thus $S_{1}+(S_{2}-_{\rm fs}S_{1})$ must be equal to $S_{2}\,$, which is impossible by Theorem 2.33. Hence again, $S_{1}+(S_{2}-_{\rm fs}S_{1})\subseteq S_{2}^{\prime}$ and therefore $S_{2}-_{\rm fs}S_{1}\subseteq S_{2}^{\prime}-_{\rm fs}S_{1}\,$. ∎